OpenStudy (anonymous):

Future Value of Wise Buying. If a person saves $63 a month by using coupons and doing comparison shopping, (a) What is the amount for year? (b) What would be the future value of this annual amount over 10 years, assuming an interest rate of 4 percent?

OpenStudy (anonymous):

would you just multiply 63 and 12 for a?

OpenStudy (anonymous):

I don't understand, I have ask my instructor for help and she wont help me so I don't know what to do so I turn to here. need some help

OpenStudy (tkhunny):

One piece at a time. Think it through and pick out all the pieces. Annual Interest Rate: 0.04 = i Monthly Interest Rate: 0.04/12 = 0.003333... = j Monthly Accumulation Factor: 1+j = 1.0033333... = r Term: (10 yrs)*(12 months / year) = 120 months Style: End of Month Accumulation Summation: \($63.00 * (r^{119} + r^{118} + .... + r^{2} + r + 1) = $63.00\cdot\dfrac{1-r^{120}}{1-r}\)